# COMMET: orders-of-magnitude speed-up in finite element method via batch-vectorized neural constitutive updates

### v1

@@ Abstract
Constitutive evaluations often dominate the computational cost of finite element (FE) simulations whenever material models are complex. Neural constitutive models (NCMs) offer a highly expressive and flexible framework for modeling complex material behavior in solid mechanics. However, their practical adoption in large-scale FE simulations remains limited due to significant computational costs, especially in repeatedly evaluating stress and stiffness. NCMs thus represent an extreme case: their large computational graphs make stress and stiffness evaluations prohibitively expensive, restricting their use to small-scale problems. In this work, we introduce COMMET, an open-source FE framework whose architecture has been redesigned from the ground up to accelerate high-cost constitutive updates. Our framework features a novel assembly algorithm that supports batched and vectorized constitutive evaluations, compute-graph-optimized derivatives that replace automatic differentiation, and distributed-memory parallelism via MPI. These advances dramatically reduce runtime, with speed-ups exceeding three orders of magnitude relative to traditional non-vectorized automatic differentiation-based implementations. While we demonstrate these gains primarily for NCMs, the same principles apply broadly wherever for-loop based assembly or constitutive updates limit performance, establishing a new standard for large-scale, high-fidelity simulations in computational mechanics.
@@ 1 Introduction
Nonetheless, ensuring physical admissibility in NCMs remains crucial. Consequently, recent work has focused on incorporating mathematical constraints into neural network architectures to enforce physical principles while preserving the networks’ expressive capacity [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Moreover, notable work has been done to obtain frameworks for appropriately training such highly-parameterized NCMs in the context of solid mechanics [17, 21, 22, 16, 23, 24]. As a result, NCMs have been successfully applied to model a broad range of material behaviors, including (anisotropic) hyperelasticity [21, 6, 9, 25, 26, 27], viscoelasticity [28, 29], plasticity [30, 31, 32, 33], generalized standard materials [18], metamaterials [34], electroelasticity [35], and thermoelasticity [36, 37].
While NCMs have demonstrated remarkable flexibility and expressivity in replicating complex material behavior, their widespread adoption is hindered by the high computational cost incurred during integration into numerical solvers, especially finite element (FE) programs. Unlike traditional constitutive models, NCMs require evaluating large computational graphs, making the calculation of stress and stiffness tensors significantly more expensive in terms of floating point operations [38], thereby causing the cost of the assembly process to overshadow that of linear solves. For example, in a standard FE calculation , comparing a Mooney-Rivlin model with an NCM trained to replicate it showed that evaluating stress and stiffness accounted for about 5% of the total computed time with the Mooney Rivlin model but rose to 54% with the NCM [39]. Even for elastic material models that require no updating of state variables, computing material behavior with an NCM can become the dominant computational bottleneck.
@@
While many studies have focused on improving FE solver performance through better CPU cache utilization and single instruction multiple data (SIMD) based vectorization [40, 41, 42, 43, 44, 45], these efforts do not address the cost of complex constitutive evaluations, assuming relatively inexpensive material models (which is not the case for NCMs). Nonetheless, the techniques developed in those works – particularly with regard to batched computations and vectorization – can inspire performance optimizations for solvers utilizing NCMs. On the other hand, several studies have employed vectorized constitutive updates for FE solves and inverse problems in solid mechanics [21, 6, 34, 35, 46, 47]. However, these efforts have primarily been implemented in non-performant languages such as Python – due to the prevalence of Python-based NCM training environments – or using JAX [48], which lacks sufficient support for scaling across multiple compute nodes. Moreover, they have not undergone a quantitative and systematic investigation into their scaling behavior or comprehensive performance analysis. Lastly, some efforts [38, 37, 7, 25] have focused on retrofitting NCMs into legacy FE solvers, including commercial software such as Abaqus [49] and Ansys [50]. While these solvers support user-defined material models and routines, their underlying architectures do not support vectorization strategies across multiple quadrature points and elements and therefore, suffer from computational bottlenecks. To fully leverage the representational power of NCMs in practical applications, it is essential that the surrounding FE solver architecture be re-imagined from the ground up.
@@
To address these challenges, we introduce a novel FE assembly algorithm that enables batched and vectorized evaluation of the constitutive model. This vectorization strategy requires simultaneous access to state variables of multiple material points, which in turn necessitates a redesign of the conventional FE element-level assembly process. To allow fine-grained control over the performance trade-offs introduced by batching, the solver architecture includes a mechanism to explicitly manage the batch size – the number of constitutive updates processed in a single vectorized operation. Typically for NCMs, the stress and stiffness are calculated by automatic differentiation, which can be slow for NCMs with large computational graphs. Therefore, we further improve FE performance by replacing automatic differentiation (AD) with compute graph optimized (CGO) implementations of NCMs. In CGO, we demonstrate that a carefully designed analytical treatment of NCMs can outperform AD by enabling efficient analytical computation of first and second derivatives. This approach significantly reduces both memory usage and computation time. Finally, we show that batch-vectorization is compatible with distributed-memory parallelism using message-passing interface (MPI) [51], effectively marrying SIMD-based parallelism within each compute node with MPI-based parallelism across multiple compute nodes. This hierarchical approach to parallelization enables highly scalable and efficient large-scale simulations.
@@
As a companion to this work, we introduce COMMET (COmputational Mechanics and Machine learning Toolbox) – an open-source software incorporating the technologies introduced in this work. Specifically, COMMET provides dedicated Python-based modules for implementing and training NCMs, as well as a C++-based FE solver with batch-vectorization, CGO, and MPI parallelism for scalable simulations using NCMs. We invite the research community to utilize and contribute to COMMET to enable broader adoption and further exploration of data-driven constitutive modeling in solid mechanics.
@@ 2.1 Finite elements for solid mechanics problems
where $K^{IJ}_{ij}$ denotes the tangent stiffness for the pair of nodes $I,J\in\{1,\,2,\,...,\,{n_{\text{nodes}}}\}$ in the spatial directions $i,j\in\{1,2,3\}$ and $r^{I}_{i}$ denotes the force residual for node $I$ in direction $i$. These entries are given by the following expressions:
@@ 2.2 Neural constitutive models
In order to satisfy the axioms of objectivity and material symmetry, the strain energy density is typically not postulated in terms of the deformation gradient $\bm{F}$ directly, but instead postulated in terms of a set of kinematic scalar values that are invariant to the choice of basis for the reference or current configuration. Examples of these scalars include the principal stretches $\lambda_{i}$, $i=1,2,3$, obtained from the spectral decomposition of $\bm{F}$ given by
$\bm{F}=\sum_{i}\lambda_{i}\bm{n}^{i}\otimes\bm{N}^{i}\,$, (8)
where $\bm{n}^{i}$ and $\bm{N}^{i}$ are the principal directions in the spatial and material configuration, respectively. Alternatively, invariants of the right Cauchy-Green tensor $\bm{C}=\bm{F}^{T}\bm{F}$, e.g.,
$I_{1}$ $=\text{tr}\left(\bm{C}\right)\,$, $I_{2}$ $=\frac{1}{2}\left[\text{tr}\left(\bm{C}\right)^{2}-\text{tr}\left(\bm{C}^{2}\right)\right]\,$, $I_{3}$ $=\text{det}\left(\bm{C}\right)\,$, (9)
$I_{4,ij}$ $=\bm{A}^{i}\cdot\bm{C}\bm{A}^{j}\,$, $I_{5,ij}$ $=\bm{A}^{i}\cdot\bm{C}^{2}\bm{A}^{j}\,$, (10)
@@
In the traditional constitutive modeling paradigm, $\mathcal{N}$ is typically a simple analytical function that is postulated by mechanicians. However, in the NCM paradigm, $\mathcal{N}$ takes the form of a neural network architecture that is agnostic to its use as a constitutive model. In accordance with machine learning (ML) terminology, we term $\mathcal{K}$ as the kinematic layer and $\mathcal{N}$ as the inner network. To ensure polyconvexity of the NCM in $\bm{F}$, certain convexity conditions are required for the inner network. These conditions are dependent on the choice of the kinematic scalars. For example, if the set of kinematic scalar inputs is $\left\{\lambda_{1},\,\lambda_{2},\,\lambda_{3},\,\lambda_{1}\lambda_{2},\,\lambda_{2}\lambda_{3},\,\lambda_{3}\lambda_{1},\,J\right\}$, then input-convexity [^1] The architecture of an input-convex neural network [55] is designed to ensure that the output is identically convex with respect to the inputs, regardless of the network’s weights. of $\mathcal{N}$ is sufficient to guarantee polyconvexity in $\bm{F}$ [12, 3]. However, if the scalar inputs are themselves polyconvex functions in $\bm{F}$ (e.g., $I_{1}$, $I_{2}$, and $I_{3}$), then one requires that $\mathcal{N}$ is convex and monotonically increasing in its inputs [6] for the NCM to be polyconvex in $\bm{F}$ overall. To this end, some inner networks that have been used to date include input convex neural networks (ICNNs) [9, 21, 55], monotonically non-decreasing input convext neural networks (MICNNs) [56], constitutive artificial neural networks (CANNs) [8, 5, 38, 57], and input convex Kolmogorov-Arnold networks (ICKANs) [6, 58]. Further details on these specific NCM-based architectures are provided in B for completeness.
@@ 3.1 Assembly vectorization and batching
An underexplored alternative is exporting trained NCM models—specifically their weights and computational graphs—from Python-based ML frameworks, and importing them into high-performance languages like C/C++, where the rest of the FE code is executed efficiently. There are two ways that this can be done: reimplement the NCM model in the C/C++ codebase by hand, or export the NCM model via a graph-compilation tool such as TorchScript [67] and launch an associated compute kernel in the C/C++ codebase. Current software environments for ML and high-performance languages like C/C++ only support the latter as hard-coding neural networks with complex architectures and millions of parameters would be too labor-intensive. However, in traditional FE assembly algorithms (see Alg.  1) that iterate over individual material points, invoking the NCM compute kernel would require launching the kernel for each material point [38], incurring substantial overhead.
To mitigate this, we leverage vectorization by processing multiple material points simultaneously through a single NCM instance. This approach reduces kernel launch overhead significantly, requiring only one invocation per batch, and thus achieving a speedup proportional to the batch size.
@@ 4.1.1 Material point benchmarks: Effect of vectorization and batching, without CGO
For each experiment, we report the absolute number of cache misses, cache misses relative to the non-vectorized baseline, absolute wall time, and speed-up in wall time compared to the non-vectorized baseline in Fig.  5. Here, the non-vectorized baseline refers to the sequential evaluation of constitutive updates for each material point in a loop, which is equivalent to a batch-vectorized computation with a batch size of 1. All material point benchmarks are performed on a workstation equipped with an AMD Ryzen 9 7950X CPU, the details of which are presented in Tab.  1 in D.
@@ 4.3 Distributed memory parallelization benchmarks
To evaluate parallelization performance of the batch-vectorized assembly algorithm (Alg.  3) and surrounding FE solver, we evaluate the strong scaling in Fig.  12, i.e. the speed-up of the wall time when using multiple processors relative to the single processor wall time, for both the single compute node and multi-compute nodes. [^2] These computational experiments were run on the Delft Blue Supercomputer [79]. See Tab.  1 in D for details of the CPUs used on these compute nodes. In a single compute node, data exchange between multiple processors is faster because all processors share the same memory. In contrast, in a multi-node setup, data must be communicated over interconnects, which can introduce latency and slow down performance.
@@ 6 Conclusion
Although our demonstrations focused on NCMs, the framework is not restricted to them: the same strategies apply wherever loop-based constitutive updates dominate runtime, from sophisticated anisotropic plasticity to multiscale homogenization. COMMET therefore lays a strong foundation not only for the practical deployment of NCMs but also for accelerating high-fidelity FE simulations more broadly across solid mechanics. Future work will target automatic batch-size tuning, support for history-dependent materials, and multiphysics extensions. Through the open-source release of COMMET, we invite the community to adopt, extend, and accelerate both neural and conventional constitutive models in computational mechanics.

### v2

@@ Abstract
Constitutive evaluations often dominate the computational cost of finite element (FE) simulations whenever material models are complex. Neural constitutive models (NCMs), i.e., neural network-based constitutive models, offer a highly expressive and flexible framework for modeling complex material behavior in solid mechanics. However, their practical adoption in large-scale FE simulations remains limited due to significant computational costs, especially in repeatedly evaluating stress and stiffness. NCMs thus represent an extreme case: their large computational graphs make stress and stiffness evaluations prohibitively expensive, restricting their use to small-scale problems. In this work, we introduce COMMET, an open-source FE framework whose architecture has been redesigned from the ground up to accelerate high-cost constitutive updates. Our framework features a novel assembly algorithm that supports batched and vectorized constitutive evaluations, compute-graph-optimized derivatives that replace automatic differentiation, and distributed-memory parallelism via MPI. These advances dramatically reduce runtime, with speed-ups exceeding three orders of magnitude relative to traditional non-vectorized automatic differentiation-based implementations. While we demonstrate these gains primarily for NCMs, the same principles apply broadly wherever for-loop based assembly or constitutive updates limit performance, establishing a new standard for large-scale, high-fidelity simulations in computational mechanics.
@@ 1 Introduction
Nonetheless, ensuring physical admissibility in NCMs remains crucial for model generalizability. This is of particular importance for the integration of NCMs in finite element (FE) solvers where non-physical material behavior typically leads to instabilities. Consequently, recent work has focused on incorporating mathematical constraints into neural network architectures to enforce physical principles as an inductive bias while preserving the networks’ expressive capacity [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Moreover, notable work has been done to obtain frameworks for appropriately training such highly-parameterized NCMs in the context of solid mechanics [17, 21, 22, 16, 23, 24]. As a result, NCMs have been successfully applied to model a broad range of material behaviors, including (anisotropic) hyperelasticity [21, 6, 9, 25, 26, 27], viscoelasticity [28, 29], plasticity [30, 31, 32, 33], generalized standard materials [18], metamaterials [34], electroelasticity [35], and thermoelasticity [36, 37].
While NCMs have demonstrated remarkable flexibility and expressivity in replicating complex material behavior, their widespread adoption is hindered by the high computational cost incurred during integration into numerical solvers, especially FE programs. Unlike traditional constitutive models, NCMs require evaluating large computational graphs, making the calculation of stress and stiffness tensors significantly more expensive in terms of floating point operations [38], thereby causing the cost of the assembly process to overshadow that of linear solves. For example, in a standard FE calculation, comparing a Mooney-Rivlin model with an NCM trained to replicate it showed that evaluating stress and stiffness accounted for about 5% of the total computed time with the Mooney Rivlin model but rose to 54% with the NCM [39]. Even for elastic material models that require no updating of state variables, computing material behavior with an NCM can become the dominant computational bottleneck.
@@
While many studies have focused on improving FE solver performance through better CPU cache utilization and single instruction multiple data (SIMD) based vectorization [40, 41, 42, 43, 44, 45], these efforts do not address the cost of complex constitutive evaluations, assuming relatively inexpensive material models (which is not the case for NCMs). Nonetheless, the techniques developed in those works – particularly with regard to batched computations and vectorization – can inspire performance optimizations for solvers utilizing NCMs. On the other hand, several studies have employed vectorized constitutive updates for FE solves and inverse problems in solid mechanics [21, 6, 34, 35, 46, 47]. However, these efforts have primarily been implemented in non-performant languages such as Python – due to the prevalence of Python-based NCM training environments – or using JAX [48], which lacks sufficient support for scaling across multiple compute nodes. Moreover, they have not undergone a quantitative and systematic investigation into their scaling behavior or comprehensive performance analysis. Lastly, some efforts [38, 37, 7, 25, 49] have focused on retrofitting NCMs into legacy FE solvers, including commercial software such as Abaqus [50] and Ansys [51]. While these solvers support user-defined material models and routines, their underlying architectures do not support vectorization strategies across multiple quadrature points and elements and therefore, suffer from computational bottlenecks. To fully leverage the representational power of NCMs in practical applications, it is essential that the surrounding FE solver architecture be re-imagined from the ground up.
@@
To address these challenges, we introduce a novel FE assembly algorithm that enables batched and vectorized evaluation of the constitutive model. This vectorization strategy requires simultaneous access to state variables of multiple material points, which in turn necessitates a redesign of the conventional FE element-level assembly process. To allow fine-grained control over the performance trade-offs introduced by batching, the solver architecture includes a mechanism to explicitly manage the batch size – the number of constitutive updates processed in a single vectorized operation. Typically for NCMs, the stress and stiffness are calculated by automatic differentiation, which can be slow for NCMs with large computational graphs. Therefore, we further improve FE performance by replacing automatic differentiation (AD) with compute graph optimized (CGO) implementations of NCMs. In CGO, we demonstrate that a carefully designed analytical treatment of NCMs can outperform AD by enabling efficient analytical computation of first and second derivatives. This approach significantly reduces both memory usage and computation time. Finally, we show that batch-vectorization is compatible with distributed-memory parallelism using message-passing interface (MPI) [52], effectively marrying SIMD-based parallelism within each compute node with MPI-based parallelism across multiple compute nodes. This hierarchical approach to parallelization enables highly scalable and efficient large-scale simulations. For the scope of this work, we focus on hyperelastic material behavior, while noting that the overall framework is agnostic to NCM architecture and can also be applied similarly to path-dependent NCMs.
@@
As a companion to this work, we introduce COMMET (COmputational Mechanics and Machine learning Toolbox) – an open-source software incorporating the technologies introduced in this work. Specifically, COMMET provides dedicated Python-based modules for implementing and training NCMs, as well as a C++-based FE solver which has been developed on top of the Deal.II library [53] with batch-vectorization, CGO, and MPI parallelism for scalable simulations using NCMs. We invite the research community to utilize and contribute to COMMET to enable broader adoption and further exploration of data-driven constitutive modeling in solid mechanics.
The remaineder of this contribution is organized as follows. Key background on finite elements and neural constitutive models is summarized in Section  2. We provide details on the global and batch-vectorized assembly algorithms as well as CGO and MPI parallelization in Section  3. The results of extensive benchmarks of the vectorization algorithms and developed solver are presented and discussed in Section  4. In Section  5, we present demonstration of the developed solver’s capabilities by simulating the passive filling of a human heart while using an NCM to model the material behavior. Finally, we provide concluding remarks in Section  6.
@@ 2.1 Finite elements for solid mechanics problems
where $K^{IJ}_{ij}$ denotes the tangent stiffness for the pair of nodes $I,J\in\{1,\,2,\,...,\,{n_{\text{nodes}}}\}$ in the spatial directions $i,j\in\{1,2,3\}$ and $r^{I}_{i}$ denotes the force residual for node $I$ in direction $i$. These entries are given by the following expressions, which result from the discretization of the weak form of the balance of equilibrium with negligible body forces in the current configuration (see, e.g. [56] Section 4.2.3):
@@ 2.2 Neural constitutive models
In order to satisfy the axioms of objectivity and material symmetry, the strain energy density is typically not postulated in terms of the deformation gradient $\bm{F}$ directly, but instead postulated in terms of a set of kinematic scalar values that are invariant to the choice of basis for the reference or current configuration. Examples of these scalars include the signed singular values of $\bm{F}$, $\nu_{i}$, $i=1,2,3$ [57, 3]. [^1] Additional requirements of $\Pi$ invariance are required for objectivity and material symmetry in this case; see [57, 3] for details. Alternatively, invariants of the right Cauchy-Green tensor $\bm{C}=\bm{F}^{T}\bm{F}$, e.g.,
$I_{1}$ $=\text{tr}\left(\bm{C}\right)\,$, $I_{2}$ $=\frac{1}{2}\left[\text{tr}\left(\bm{C}\right)^{2}-\text{tr}\left(\bm{C}^{2}\right)\right]\,$, $I_{3}$ $=\text{det}\left(\bm{C}\right)\,$, $I_{4,ij}$ $=\bm{A}^{i}\cdot\bm{C}\bm{A}^{j}\,$, (8)
@@
In the traditional constitutive modeling paradigm, $\mathcal{N}$ is typically a simple analytical function that is postulated by mechanicians. However, in the NCM paradigm, $\mathcal{N}$ takes the form of a neural network architecture that is agnostic to its use as a constitutive model. In this sense, $\mathcal{N}$ is simply a highly expressive ansatz for a constitutive equation that is inspired by machine learning. In accordance with machine learning (ML) terminology, we term $\mathcal{K}$ as the kinematic layer and $\mathcal{N}$ as the inner network. To ensure polyconvexity of the NCM in $\bm{F}$, certain convexity conditions are required for the inner network. These conditions are dependent on the choice of the kinematic scalars. For example, if the set of kinematic scalar inputs is $\left\{\nu_{1},\,\nu_{2},\,\nu_{3},\,\nu_{1}\nu_{2},\,\nu_{2}\nu_{3},\,\nu_{3}\nu_{1},J\right\}$, then input-convexity [^2] The architecture of an input-convex neural network [58] is designed to ensure that the output is identically convex with respect to the inputs, regardless of the network’s weights. of $\mathcal{N}$ is sufficient to guarantee polyconvexity in $\bm{F}$ [57, 12, 3]. However, if the scalar inputs are themselves polyconvex functions in $\bm{F}$ (e.g., $I_{1}$, $I_{2}$, and $I_{3}$), then one requires that $\mathcal{N}$ is convex and monotonically increasing in its inputs [6] for the NCM to be polyconvex in $\bm{F}$ overall. To this end, some inner networks that have been used to date include input convex neural networks (ICNNs) [9, 21, 58], monotonically non-decreasing input convext neural networks (MICNNs) [59], constitutive artificial neural networks (CANNs) [8, 5, 38, 60], and input convex Kolmogorov-Arnold networks (ICKANs) [6, 61]. Further details on these specific NCM-based architectures are provided in B for completeness.
@@ 3.1 Assembly vectorization and batching
An underexplored alternative is exporting trained NCM models—specifically their weights and computational graphs—from Python-based ML frameworks, and importing them into high-performance languages like C/C++, where the rest of the FE code is executed efficiently. There are two ways that this can be done: reimplement the NCM model in the C/C++ codebase by hand, or export the NCM model via a graph-compilation tool such as TorchScript [70]. Current software environments for ML and high-performance languages like C/C++ only support the latter as hard-coding neural networks with complex architectures and large numbers of parameters would be too labor-intensive.
Each call that is made to the graph-compiled model requires several overhead operations, e.g. the data type, layout, and contiguity of the input tensors is checked, reference counters are updated, memory is allocated, etc. The cost of these operations is known as the “compute kernel launch overhead”. This cost is distinct from the cost of instantiating an NCM and loading its weights from memory, i.e. even for a single NCM instance the compute kernel launch overhead is incurred for each call made to that instance. In traditional FE assembly algorithms (see Alg.  1), the compute kernel is launched for every material point [38], and so the incurred compute kernel launch overhead is multiplied by the number of material points.
To mitigate this, we leverage vectorization by processing multiple material points simultaneously. This approach reduces the incurred kernel launch overhead significantly, requiring only one invocation per batch, and thus achieving a speed-up proportional to the batch size.
@@
Remark.
The necessity of instructions being identical across all data poses a challenge for constitutive updates that may have control flow which differs from one material point to the next, e.g. updates that require local Newton-Raphson iterations for the so-called return mapping algorithm in plasticity. A possible solution in this case is to continue iterating across the whole batch until the state variables for all material points have converged. We note, however, that many path dependent NCMs that have been presented in the literature do not require solving of nonlinear equations for the evolution state variables through such iterative schemes [28, 71, 30, 31], and so the proposed global and batch-vectorized algorithms can be applied as is.
@@ 4.1.1 Material point benchmarks: Effect of vectorization and batching, without CGO
For each experiment, we report the absolute number of cache misses, cache misses relative to the non-vectorized baseline, absolute wall time, and speed-up in wall time compared to the non-vectorized baseline in Fig.  5. Here, the non-vectorized baseline refers to the sequential evaluation of constitutive updates for each material point in a loop, which is equivalent to a batch-vectorized computation with a batch size of 1. All material point benchmarks are performed on a workstation equipped with an AMD Ryzen 9 7950X CPU, the details of which are presented in Tab.  2 in D.
@@ 4.3 Distributed memory parallelization benchmarks
To evaluate parallelization performance of the batch-vectorized assembly algorithm (Alg.  3) and surrounding FE solver, we evaluate the strong scaling in Fig.  12, i.e. the speed-up of the wall time when using multiple processors relative to the single processor wall time, for both the single compute node and multi-compute nodes. [^3] These computational experiments were run on the Delft Blue Supercomputer [82]. See Tab.  2 in D for details of the CPUs used on these compute nodes. In a single compute node, data exchange between multiple processors is faster because all processors share the same memory. In contrast, in a multi-node setup, data must be communicated over interconnects, which can introduce latency and slow down performance.
@@ 6 Conclusion
Although our demonstrations focused on NCMs, the framework is not restricted to them: the same strategies apply wherever loop-based constitutive updates dominate runtime, from sophisticated anisotropic plasticity to multiscale homogenization. COMMET therefore lays a strong foundation not only for the practical deployment of NCMs but also for accelerating high-fidelity FE simulations more broadly across solid mechanics. Future work will target automatic batch-size tuning, support for history-dependent materials, multiphysics extensions, and GPU implementation. Through the open-source release of COMMET, we invite the community to adopt, extend, and accelerate both neural and conventional constitutive models in computational mechanics.